OpenStudy (anonymous):

the product of 4.6 and a number is 34.2 more than the number

OpenStudy (anonymous):

The product of 4.6 and a number: 4.6 times x: 4.6x is = 34.2 more than the number: 34.2 + x

OpenStudy (anonymous):

how do i do it do i subtract 34.2 from each side?

OpenStudy (anonymous):

\[4.6x = 34.2+x\] \[4.6x-x=34.2\] \[3.6x = 34.2\] \[x=\frac{34.2}{3.6}\]

OpenStudy (anonymous):

how did u get 3.6 im sorry im just really confuzed

OpenStudy (anonymous):

4.6x-1x = x(4.6-1) just like if you had 3 x's and took away 1 you would have 2

OpenStudy (anonymous):

ohs ok im sorry im bad at math

OpenStudy (anonymous):

so now i just have to divide 34.2 by 3.6?

OpenStudy (anonymous):

9.5?

OpenStudy (anonymous):

yes:) And you aren't bad at math, you're working on it. yes 9.5!

OpenStudy (anonymous):

bob bought a popccorn a hotdog and a soda at the movies for 8.83 popcorn costs one dollar more than a hotdog a soda costs 0.42 cents less than a hotdog how much is the item

OpenStudy (anonymous):

Well, what item? The problem breaks down as: P=popcorn H=hotdog S = soda P + H +S = 8.83 P+1=H S-0.42 = H

OpenStudy (anonymous):

the first 1 is trying to figure out how much the hot dog is

OpenStudy (anonymous):

Ok, So we need to somehow substitute values of H and money into P and S in that first equation.

OpenStudy (anonymous):

haha phs was my highschool initials but im in college now but im doing prealgebra

OpenStudy (anonymous):

P + H +S = 8.83 P+1=H --> P = H-1 S-0.42 = H ---> S=H+0.42 so P + H +S = 8.83 becomes H-1 + H +H+0.42 = 8.83 becomes 3H-1 +0.42 = 8.83 3H = 8.83+1-0.42 3H = 9.83-0.42 = 9.42 H = 9.42/3

OpenStudy (anonymous):

3.14?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

expensive hotdog

OpenStudy (anonymous):

it said please try again

OpenStudy (anonymous):

?

OpenStudy (anonymous):

3H = 9.83-0.42 = 9.42 WRONG = 9.41 H = 9.41/3 I don't know if this will make a difference though. H=3.13 and 2/3

OpenStudy (anonymous):

it said it was 2.75

OpenStudy (anonymous):

cause i tried that part to many times

OpenStudy (anonymous):

its just the first one thats hardd then the other 2 are easy i got a similar question cause o missed the first 1

OpenStudy (anonymous):

9.86 popcorn hot dog soda popcorn is 1 dollar more than a hot dog soda costs 0.89 cents less than a hotdog